AM 4 -Esercitazione 2 A.A.2003-2004

Prof. Luigi Chierchia, Dott. Laura Di Gregorio 2 ottobre 2003

Esercizi

Calcolare i seguenti integrali indefiniti:

1. ∫ \frac{dx}{(x + 1)^{2}\sqrt{x^{2} + 2x + 2}}
2. ∫ \frac{dx}{(x − 1)\sqrt{x^{2} − 3x + 2}}

3. ∫ \sqrt{3 − 2x − x^{2}} dx

4. ∫ \sqrt{x^{2} − 4} dx

5. ∫ x\sqrt{x^{2} + x + 1} dx